# Bifröst: 3D-Aware Image compositing with Language Instructions

### v1

@@ Appendix C Evaluation Details
TF-ICON Evaluation. When testing on TF-ICON, we employ the parameter set in "same domain" mode, as suggested by the official implementation. It also requires a text prompt as an additional input, so we apply BLIP2 (Li et al., 2023a), a state-of-the-art vision-language model to generate captions for the images. Moreover, the captions for the DreamBooth (Ruiz et al., 2023) test set are manually refined to improve the performance.
@@ Appendix F Human Evaluation Interface
    We chose 30 images from DreamBooth (Ruiz et al., 2023) test set, and for each image, we then generated 3 variants with different backgrounds using each models, respectively. Overall, there were 30 original images and 90 generated variants in total.
@@
    We gathered 30 volunteers from various backgrounds. We prepare detailed regulations and templates for annotators to rate the images for scores of 1 to 5 from three perspectives: “Fidelity”, “quality”, and “diversity”. “Fidelity” measures the ability of ID preserving. “Quality” counts for whether the generated image is harmonized without considering fidelity. As we do not encourage “copy-paste” style generation, we use “diversity” to measure the differences among the generated proposals.

### v2

@@ Appendix F Human Evaluation Interface
    We chose 30 images from DreamBooth (Ruiz et al., 2023) test set, and for each image, we then generated 3 variants with different backgrounds using each model, respectively. Overall, there were 30 original images and 90 generated variants in total.
@@
    We recruited 30 volunteers from diverse backgrounds and provided detailed guidelines and templates for evaluation. Annotators rated the images on a scale of 1 to 5 across four criteria: “Fidelity”, “Quality”, “Diversity”, and “3D Awareness”. “Fidelity” evaluates identity preservation, while “Quality” assesses visual harmony, independent of fidelity. “Diversity” measures variation among generated proposals to discourage “copy-paste” style outputs. “3D Awareness” evaluates the object’s ability to handle spatial relationships (e.g., occlusion) seamlessly.